In Problem of the solution as t- Oo 16,solve the given initial value problem. Describe the behavior 2 16. х — 5 х, х(0) 4
Answer #1

slution X = (-2 ) 1-5 4 Let n = ) 2 1-5 4/22/- (-50g +472/ » Nů = -22, the and a = -5 01 +4012 dy = -24 the due = -5h, the >Putting value of ns (t) and dry in any = -21, the , we get my as z (t) = dry +226 = 39 34 - Get + zgest t2G et 374) = 5934 +(:es and E0 = %= % As ko, nito. Use initial condition > N; (0) = 1 and 2(o) = 3 1 = Gtel and 3=5966 (1 =1) G= ² subtract -2=-